dynamics:
A 50 kg person is walking horizontally with a constant acceleration of 0.25 m/s squared inside an elevator. The elevator is also accelerating downward at a rate of 1 m/s squared.
a) find the force of friction between his shoe and the floor of the elevator?
b) what minimum coefficient of friction between the shoe and the floor it would make this scenario possible?
1
Expert's answer
2020-07-13T11:44:59-0400

a)


"F_{fr}=ma_2=50\cdot0.25=12.5N"


b)


"N=mg-ma_1=50\cdot9.81-50\cdot 1=441N"


"F_{fr}=\mu N\to \mu=\frac{F_{fr}}{\mu}=\frac{12.5}{441}=0.028"